Ask your own question, for FREE!
Mathematics 13 Online
OpenStudy (anonymous):

Tasha’s bowl contains 4 red, 6 green and 3 brown candies. She randomly chooses and keeps 3 candies from the bowl. What is the probability she will choose all brown?

OpenStudy (anonymous):

Here is what I did, in case anybody needs help in the future. I did the following: Total outcomes; 4 + 6 + 3 = 13 First draw: 3/13 Second draw: 2/12 Third draw: 1/11 3/13 * 2/12 * 1/11 = 1/286

OpenStudy (bahrom7893):

3/13 * 2/12 * 1/11

OpenStudy (bahrom7893):

that looks correct

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!